# E-TRIALS: Empowering Data-Driven Decisions to Enhance Computer-Based Learning Platforms

### v1

@@ 2.1.3 Deploying an Experiment
  • Automatic Deployment: Studies categorized as single support or support comparison are automatically deployed. These studies are added to ASSISTments’ content library, where teachers can assign them to students. Randomization occurs at the student level within the platform.
  • Self-Deployment: Problem varied studies require self-deployment because researchers need to recruit their own participation pool. Researchers collaborate with students, teachers, or schools to assign content. Randomization for these studies can occur at various levels, include student, teacher, class, or school.
@@ 2.1.4 Completion of Experiment
Researchers will receive eight distinct data files, each containing specific insights into student interactions, learning outcomes, and experimental conditions. These files provide a comprehensive view of the study’s results, ranging from click stream data to problem difficulty and demographic context.
@@
    Assignment Logs: This file summarizes assignment-level data. Each row includes information on total time spent on the assignment, the total number of attempts or hints, and the cumulative discrete score for the assignment.
@@
    Prior Logs: This files provides class- and school-level data related to students’ prior learning. For the class-level data, it includes the curriculum used, the class grade level, the number of students in the class, and the class’s average performance. For the school-level data, we get this data from the National Center for Education Statistics (NCES) [Hussar and Bailey (2011)] based on the teacher’s school address. It includes demographic data, such as percentage of students receiving free or reduced lunch, ethnicity and gender distribution, and the school’s location (rural, urban, or suburban). The school_id provided in this file can be used to access additional school-level, or even district-level data through NCES or the Education Data Explorer by the Urban Institute. If a teacher does not provide their school address, however, it is not feasible to link the data for their class to school-level data.
@@
    Assignment Settings: The file details the settings applied to the assignment. Assignment settings be setting a time limit or enabling the redo feature. The redo feature allows students to attempt a similar, but not the same problem again if a student fails to achieve a perfect score on the initial problem.
@@
These eight files are available only for studies conducted with participants from ASSISTments. Studies involving external participants (e.g. college students recruited for problem-varied studies) will not include data on students’ prior and post performance in ASSISTments. To view an example of the eight data files, click here.
@@ 3 Example Studies
3.1 The Three Experiments
One unique advantage of E-TRIALS is the consistent format and structure of the eight data files provided to researchers after the completion of an experiment. This consistency can help researchers reuse the same code across multiple experiments. To demonstrate this advantage, we selected three support comparison studies within E-TRIALS and applied the same code to evaluate all three experiments. A link to the code is provided in the Data Cleaning & Preparation section.
The three experiments are summarized below:
@@ 3.1 The Three Experiments
    Experiment 14 (similar vs. different example feedback): This experiment explored math learning and persistence by comparing two feedback conditions:
@@
    Students were randomly assigned to one of the two conditions. Figure 2 illustrates the differences between these feedback types.
    Figure 2: The support messages students received in both conditions
@@
    Experiment 15 (drawing vs. imagining the problem vs. image provided): This study investigate the impact of visualization techniques. Students were randomly assigned to one of three conditions:
    • Drawing on Paper: Students were asked to physically draw the image in the problem.
    • Mental Visualization: Students visualized the image mentally without drawing.
    • Provided Drawing: Students were given a completed drawing of the image.
    Figure 3 highlights the differences between the three conditions.
    Figure 3: The support messages students received in each condition
  3. 3.
    Experiment 16 (reflect vs. regular hint): This experiment examined the effectiveness of encouraging students to reflect on their confusion before requesting an additional hint.
@@
    Students were randomly assigned to one of two conditions. Figure 4 displays the reflection prompt.
    Figure 4: The support messages students received in both conditions
@@ 3.2 Data Cleaning & Preparation
To ensure we analyze the experiments correctly, several filtering steps were applied. First, we removed ASSISTments’ default condition, best-so-far because it provides students with different support messages using different format. Meaning, one student could receive a support message in a text format, whereas another student could receive it in a video. Due to the inconsistency, we decided to remove the condition from our analysis.
Next, we removed students who did not request at least one hint during the assignment, as hint interaction was required to determine the effect of the experiment. This filtering step resulted in the removal of 400 students in experiment 14, 109 students in experiment 15, and 477 students in experiment 16. Additionally, sub-problems were treated as independent problems. The updated summary of the dataset after applying these filters is presented in Table 2.
To assess the extend to student participation, we calculated the attrition rate. We calculated the attrition rate by using the proportion of missing values in the post and prior continuous score variables that’s in the Same Skill Prior & Post Logs tables relative to the total number of observations [Schafer and Graham (2002)]. The attrition rates for the three experiments were significantly low, indicating that the vast majority of participants competed the study.
@@ 3.3 Feature Selection in Covariate Adjustment
It is essential for students to have comparable prior knowledge and learning motivation. To minimize the influence of confounding factors, we aim to enhance comparability through covariate adjustment. Fortunately, our dataset contains a large number of covariates, which we leverage to account for variation in students’ baseline knowledge.
First, we utilized students’ prior and post continuous scores from assignments associated with the same skill code as the experimental problems. This data was drawn from the Same Skill Prior Logs and Same Skill Post Logs. Second, we analyzed the Prior Logs table to identify covariates potentially correlated with student performance. Although numerous covariates are available, feature selection is necessary to ensure computational efficiency, particularly for OLS regression. For Random Forest models, however, feature selection is handled inherently, as the algorithm randomly selects subsets of features during the splitting process.
We began by subjectively selecting a set of covariates. Next, we calculated the linear correlations between these covariates and student performance, visualizing the results in Figure 5. While focusing solely on linear correlation is not the most comprehensive approach, as it overlooks non-linear relationships, we chose this measure due to its simplicity, ease of application, and suitability for OLS regression. [^1] OLS regression and LOOP, based on Random Forest, will be discussed in detail in the next section.
Figure 5: Feature Selection
Based on Figure 5, we excluded covariates in the plot that appeared close to white, as they indicated weak correlations. This process allowed us to identify covariates with meaningful correlations to the outcomes through a data-driven approach.
Finally, the assignment-level dataset was combined with the prior and post logs to create the final dataset used for analysis.
@@ 4 Method
In this section, we discuss the statistical methods employed to assess the impact of different types of tutoring support on students’ proximal and distal performance. We also provide the rationale for selecting these methods, emphasizing their respective advantages and limitations. The methods are presented in formal mathematical terms for clarity and rigor. The two outcome variables we explored are distal and proximal learning. Distal learning refers to the knowledge students retain after the experiment. It is measured by the average score students received on problem sets post-experiment, which assess the same skills covered during the experiment. Proximal learning refers to the knowledge students gain during the experiment. It is measured by the average continuous score students received during the experiment problem set.
@@
Assume each participant has two nonrandom potential outcomes– treatment potential outcomes $t_{i}$ and control potential outcomes $c_{i}$. We can write observed outcomes $Y_{i}$ as
@@
which serves as our vital interested parameter to be estimated.
Below we introduce LOOP estimator $\hat{\tau}_{LOOP}$ along with the other two common ATE estimators: Student’s t-test estimator $\hat{\tau}_{t-test}$ and Ordinary Least Squares (OLS) regression estimator $\hat{\tau}_{OLS}$. We also explore their strengths and limitations.
@@ 4.1 Student’s t-test
$\hat{\tau}_{t-test}$ represents the difference of the average of the observed treatment outcomes and the average of the observed control outcomes.
@@
where $n$ be the number of participants in treatment arm. It provides an unbiased estimate of the average treatment effect with $0<n<N$.
4.2 OLS Regression
Outcomes $Y_{i}$ and $q$-dimensional baseline covariates $Z_{i}$ are available in E-Trials, enabling the implementation of covariate adjustments for estimating the ATE. OLS regression estimators $\hat{\tau}_{OLS}$ naturally arise when considering covariate adjustment.
$Y_{i}=\tau_{OLS}T_{i}+\beta X_{i}+\epsilon_{i}$ (6)
where $\hat{\tau}_{OLS}$ can serve as one estimator for ATE $\tau$.
The estimator $\hat{\tau}_{\text{OLS}}$ leverages information from both treatment assignment $T_{i}$ and $q$-dimensional baseline covariates $Z_{i}$, providing a more precise estimate compared to $\hat{\tau}_{\text{t-test}}$. However, $\hat{\tau}_{\text{OLS}}$ has two notable limitations: (a) it may be biased since the assumption $X_{i}\perp T_{i}$ is violated, and (b) it fails to account for potential non-linear relationships between $Y_{i}$ and $X_{i}$.
@@ 4.3 LOOP
To address the limitations of the OLS estimator, we propose using LOOP to estimate $\tau$. $\hat{\tau}_{LOOP}$ offers an unbiased estimate of ATE while incorporating covariate adjustments. We present the steps of the LOOP methodology. A key quantity, $m_{i}$, is defined for each participant as follows:
@@
$=\tau_{i}$
Additionally, we can demonstrate that $\hat{\tau}_{i}$ exhibiting reduced variance, provided that $\hat{m}_{i}$ closely approximates $m_{i}$. The variances can be calculated as follows:
@@
When it comes to methods for imputing potential outcomes, simply ignoring covariates during imputation makes the LOOP estimator equivalent to the student’s t-test estimator. However, the LOOP estimator outperforms $\hat{\tau}_{\text{SD}}$ as long as the imputation of potential outcomes improves beyond the baseline mean imputation approach. In the work by Wu and Gagnon-Bartsch [Wu and Gagnon-Bartsch (2018)], Random Forest (RF) is recommended for imputing $t_{i}$ and $c_{i}$. This is because RF, as an enhancement over individual decision trees, not only improves accuracy but also provides automatic variable selection, resulting in more computational efficiency.
In this paper, we calculate the t-test estimator $\hat{\tau}_{\text{t-test}}$, the OLS regression estimator $\hat{\tau}_{\text{OLS}}$, and the LOOP estimator $\hat{\tau}_{\text{LOOP}}$ to demonstrate that the LOOP estimator is an unbiased estimator of the ATE and achieves greater precision by adjusting for baseline covariates. Furthermore, when using the LOOP estimator with Random Forest (RF), we can identify non-parametric relationships within the data that the OLS estimator fails to capture.

### v2

@@ 2.1.3 Deploying an Experiment
  • Automatic Deployment: Studies categorized as single support or support comparison are automatically deployed. These studies are added to ASSISTments’ content library, where teachers can assign them to students.
  • Self-Deployment: Problem varied studies require self-deployment because researchers need to recruit their own participation pool. Researchers collaborate with students, teachers, or schools to assign content.
One prerequisite to deploying a study, regardless of the type, is that the researcher would need to get IRB approval from their institution. Once they attain IRB approval, then the study can be deployed and participants can begin interacting with the intervention.
@@ 2.1.4 Completion of Experiment
Before receiving the intervention data, researchers are required to pre-register their study at OSF.io. Once the study has pre-registration is complete, researchers receive eight distinct data files, each containing specific insights into student interactions, learning outcomes, and experimental conditions. These files provide a comprehensive view of the study’s results, ranging from click stream data to problem difficulty and demographic context.
@@
    Assignment Logs: This file summarizes assignment-level data. Each row includes information on the total time spent on the assignment, the total number of attempts or hints, and the cumulative discrete score for the assignment.
@@
    Prior Logs: This file provides class- and school-level data related to students’ prior learning. For the class-level data, it includes the curriculum used, the class grade level, the number of students in the class, and the class’s average performance. School-level data can be provided if a teacher provides their school address in their profile. Then the school address is used to retrieve school-level data from the National Center for Education Statistics (NCES) [Hussar and Bailey (2011)]. School-level data includes demographic data, such as percentage of students receiving free or reduced lunch, ethnicity, and gender distribution, and the school’s location (rural, urban, or suburban). The school_id provided in this file can be used to access additional school-level, or even district-level data through NCES or the Education Data Explorer by the Urban Institute. If a teacher does not provide their school address, however, it is not feasible to link the data for their class to school-level data.
@@
    Assignment Settings: The file details the settings applied to the assignment. Some are setting a time limit or enabling the redo feature. The redo feature allows students to attempt a similar, but not the same problem again if a student fails to achieve a perfect score on the initial problem.
@@
These eight files are available only for studies conducted with participants from ASSISTments. Studies involving external participants (e.g., college students recruited for problem-varied studies) will not include data on students’ prior and post-performance in ASSISTments. An example of the eight data files can be viewed here.
@@ 3 Example Studies
3.1 The Two Experiments
In this section, we introduce two E-TRIALS support comparison studies that we’ll use to showcase an example of analyzing E-TRIALS data to determine the efficacy of the studies. We’ll explain what each experiment explored and their conditions. A link to the code is provided in the Data Cleaning & Preparation sub-section. Data for the two experiments can be provided upon request.
The two experiments are summarized below:
  • Experiment 1 (similar vs. different example feedback): This experiment explored math learning and persistence by comparing two feedback conditions:
@@ 3.1 The Two Experiments
    Students were randomly assigned to one of the two conditions. Figure 2 illustrates the differences between these feedback types. The Similar figure informs students in the first hint that the subsequent hint will be a new problem completely worked out, and they should reference the steps to solve the original problem. The new problem that appears in the subsequent hint is similar to the original problem, whereas in the Different figure, the new problem that appears in the second hint is different than the original problem. Click here to view the figures in higher resolution.
    Figure 2: The support messages students received in both conditions. The left figure is the Similar condition, and the right figure is the Different condition.
  • Experiment 2 (reflect vs. regular hint): This experiment examined the effectiveness of encouraging students to reflect on their confusion before requesting an additional hint.
    1. 1.
@@
    Students were randomly assigned to one of two conditions. Figure 3 displays the reflection prompt. The first hint in the Reflect figure prompts students to reflect on their knowledge, and to write down on a piece of paper the parts of the problem they understand and the parts they find confusing. In the Control figure, students receive hints as they normally would in ASSISTments. Click here to view the figures in higher resolution.
    Figure 3: The support messages students received in both conditions. The left figure is the Reflect condition, and the right figure is the Control condition.
3.2 Identifying Target Populations while Avoiding Selection Bias
Both experiments under consideration contrasted student supports—just-in-time homework help that struggling students can receive upon request. In both cases, our interest is in those students who requested help from the tutoring system, so we conducted our analysis on the subset of the data consisting of students who, at some point during the experiment, requested such help.
In general, depending on the design of the A/B testing platform and other specifics, this type of subsetting, as innocuous as it may seem, can lead to selection bias. For instance, say that students who request support are penalized for doing so, and the extent of the penalty is different for the two conditions in a study. For the sake of argument, say the penalty is greater in the treatment condition than in control. In that case, we might expect that some relatively strong students (those who are deciding between requesting support and trying to figure the problems out on their own) would be dissuaded from requesting hints if assigned to the treatment condition but not if assigned to control. If they are then dropped from the analysis sample, the average learning outcome in the treatment group will be biased downward. This is an example of post-randomization selection bias [Li et al. (2024)] that arises when researchers analyze subsets of their data that might have been affected by treatment assignment.
Fortunately, the current version of the E-TRIALS platform has been designed to avoid this problem (though it affected earlier designs—for instance, leading [Patikorn and Heffernan (2020)] to avoid subsetting). In order to request support, students must first press a button labeled “Get Help.” The appearance of this button does not depend on the type of supports that are available—in particular, it is identical for both treatment arms in our two example studies. Hence, before the first time during the experiment a student pressed “Get Help,” they had no way of knowing their treatment randomization. Therefore, the excluded subset of students who never pressed “Get Help” cannot be affected by treatment randomization—the subset would have included the same students under any permutation of the treatment assignment, and its exclusion does not cause selection bias.
@@ 3.3 Data Cleaning & Preparation
To ensure we analyze the experiments correctly, several filtering steps were applied. First, we removed ASSISTments’ default condition, best-so-far. Once again, this condition represents the highest-performing student support currently in ASSISTments for that problem. Meaning, one student could receive a support message in a text format, whereas another student could receive it in a video. Due to the inconsistency, we decided to remove the condition from our analysis.
Next, we removed students who did not request at least one hint during the assignment, as hint interaction was required to determine the effect of the experiment. This filtering step resulted in the removal of 400 students from Experiment 1 and 477 students from Experiment 2.
Lastly, several assignments in ASSISTments contain sub-problems. A sub-problem is linked to a parent problem and builds on the same scenario or context. Sub-problems often guide students to think conceptually or to apply the result from the parent-problem to solve the sub-problem. For example, a parent problem can set out the scenario or context and then ask the student to calculate the commission from the sale. Then, the sub-problem can ask the student to use the results from the parent-problem to determine how much money the store receives after the commission is deducted. With that, it’s important to note that none of the sub-problems in the two experiments were conceptual, where a student was prompted to provide a reasoning for their answer to the parent problem. Since none of the sub-problems were conceptual, we treated them as independent problems. The updated summary of the dataset after applying these filters is presented in Table 2.
To assess student participation across the study, we calculated the attrition rate, which refers to the loss of participants over the course of a study [Clearinghouse (2015)]. Attrition occurs when students who were initially randomly assigned to a condition do not provide outcome data. In our case, we measured the attrition rate using the proportion of missing values in the pre- and post-intervention continuous score variables from Same Skill Prior & Post Logs tables relative to the total number of observations [Schafer and Graham (2002)]. The attrition rates for the two experiments were low, which means the vast majority of participants completed the study and had outcome data.
@@ 3.4 Feature Selection in Covariate Adjustment
It is essential for students to have comparable prior knowledge and motivation to learn. The RCTs within ASSISTments use Bernoulli randomization, so students in the treatment and control arms should be comparable on average. However, in practice, slight imbalances in the students’ baseline characteristics may still exist. To assess the comparability of students in the treatment and control arms of our two selected experiments, we performed a classification permutation test (CPT). This test evaluates whether the randomization procedure adequately mitigates potential confounders arising from differences in prior knowledge and learning motivation. Table 3 shows that while there is no strong evidence of imbalance between the treatment and control groups, some degree of covariate difference remains, suggesting that covariate adjustment is still necessary.
Table 3: Covariate Balance Check
Experiments p-value
1 0.196
2 0.124
To minimize the influence of confounding factors, our aim is to enhance comparability through covariate adjustment. Fortunately, our data set contains a large number of covariates, which we leverage to account for variations in the students’ baseline knowledge.
First, we utilized students’ prior and post-continuous scores from assignments associated with the same skill code as the experimental problems. These data were drawn from the Same Skill Prior Logs and Same Skill Post Logs. Second, we analyzed the Prior Logs table to identify covariates that could be correlated with student performance. Although numerous covariates are available, feature selection is necessary to ensure computational efficiency and reduce multicollinearity, particularly for regression. However, in LOOP, which is based on Random Forest (RF) models, feature selection is handled inherently. At each split in a tree, the algorithm randomly selects a subset of features and chooses the one that best splits the data based on a predefined criterion, making feature selection a local and adaptive process. 111regression and LOOP, based on Random Forest, will be discussed in detail in the next section.
We began by selecting a set of covariates based on their variability and pairwise correlations. Specifically, we removed features with near-zero variance, as they carry little informational value, and eliminated highly correlated variables to reduce multicollinearity. Although relying solely on linear correlation does not capture nonlinear dependencies, we adopted this approach for its simplicity, ease of implementation, and compatibility with linear mixed-effects models such as lmer. After applying feature selection procedures, we retained a set of covariates that exhibit sufficient variability and are not highly correlated with each other. These variables were selected to ensure meaningful input into the linear mixed-effects modeling process, while reducing redundancy and potential multicollinearity. The final set includes student-level, class-level, and performance-related variables, as listed in Table 4.
Table 4: Selected Variables After Feature Selection
Variable Name Description
similar Binary indicator of treatment or control condition
class_xid Unique identifier for each class
problems_assigned Number of problems assigned
student_prior_median_first_repsonse_time_task Median first response time (prior assignments)
student_prior_median_time_on_task Median time on task (prior assignments)
student_prior_problem_sets_completed Number of problem sets completed previously
student_prior_average_attempt_count Average number of attempts (prior assignments)
student_prior_viewed_assignment_report_percent Percent of assignment reports viewed
prior_avg_cont_score Average pre-test continuous score
Finally, the assignment-level dataset was combined with the prior and post-draft logs to create the final dataset used for analysis.
@@ 4 Method
In this section, we discuss the statistical methods employed to assess the impact of different types of tutoring support on students’ proximal and distal performance. The two outcome variables that we explored are distal and proximal learning. Distal learning ($avg\_score$) refers to the knowledge that students retain after the experiment. It is measured by the average score students received on problem sets post-experiment, which assesses the same skills covered during the experiment. Proximal learning ($post\_avg\_score$) refers to the knowledge that students gain during the experiment. It is measured by the average continuous score that the students received during the set of experiment problems. We also provide the rationale for selecting these methods, emphasizing their respective advantages and limitations. The methods are presented in formal mathematical terms for clarity and rigor.
@@
Assume each participant has two nonrandom potential outcomes– treatment potential outcomes $t_{i}$ and control potential outcomes $c_{i}$ [Rubin (1974), Splawa-Neyman et al. (1990)]. We can write observed outcomes $Y_{i}$ as
@@
which is the main parameter we’re interested in estimating.
Below we introduce LOOP estimator $\hat{\tau}_{LOOP}$ along with the other two common ATE estimators: Student’s t-test estimator $\hat{\tau}_{t-test}$ and Regression estimator $\hat{\tau}_{Reg}$. We also explore their strengths and limitations.
@@ 4.1 Student’s t-test
$\hat{\tau}_{t-test}$ represents the difference between the average of the observed treatment outcomes and the average of the observed control outcomes.
@@
where $n$ is the number of participants in the treatment arm. It provides an unbiased estimate of the average treatment effect with $0<n<N$.
4.2 Regression
Outcomes $Y_{i}$ and $q$-dimensional baseline covariates $X_{i}$ are available in E-Trials, enabling the implementation of covariate adjustments for estimating the ATE. Regression estimators $\hat{\tau}_{Reg}$ naturally arise when considering covariate adjustment.
$Y_{ij}=\tau_{\text{Reg}}\cdot T_{ij}+\mathbf{X}_{ij}^{\prime}\boldsymbol{\beta}+u_{j}+\epsilon_{ij}$ (6)
where $Y_{ij}$ is the outcome for student $i$ in class $j$, $T_{ij}$ is a binary treatment indicator, $\mathbf{X}_{ij}$ is a vector of observed covariates, $u_{j}\sim\mathcal{N}(0,\sigma_{u}^{2})$ is a class-level random intercept, and $\epsilon_{ij}\sim\mathcal{N}(0,\sigma^{2})$ is the individual error term. Under the assumptions of no unmeasured confounding given $\mathbf{X}_{ij}$, correct model specification, and no interference, the coefficient $\hat{\tau}_{\text{Reg}}$ provides an unbiased estimate of ATE, $\tau$.
By adjusting for treatment $T_{ij}$ and covariates $\mathbf{X}_{ij}$, $\hat{\tau}_{\text{Reg}}$ typically improves precision over the unadjusted estimator $\hat{\tau}_{\text{t-test}}$. However, it may be biased if $\mathbf{X}_{ij}\not\perp T_{ij}$ due to imperfect randomization, and it assumes linearity between covariates and outcomes, potentially overlooking nonlinear effects.
@@ 4.3 LOOP
To address the limitations of the Regression estimator, we propose using LOOP to estimate $\tau$. $\hat{\tau}_{LOOP}$ offers an unbiased estimate of ATE while incorporating covariate adjustments [Wu and Gagnon-Bartsch (2018)]. We present the steps of the LOOP methodology. A key quantity, $m_{i}$, is defined for each participant as follows:
@@
Also we have
$Y_{i}U_{i}=\begin{cases}\frac{t_{i}}{p_{i}},&T_{i}=1,\\ -\frac{c_{i}}{1-p_{i}},&T_{i}=0.\end{cases}$ (9)
@@
$=\frac{t_{i}}{p_{i}}P(T_{i}=1)+\frac{-c_{i}}{1-p_{i}}P(T_{i}=0)+E(\hat{m_{i}})\cdot 0$
$=t_{i}-c_{i}+0$
$=\tau_{i}$ (10)
Additionally, we can demonstrate that $\hat{\tau}_{i}$ has reduced variance compared to $\hat{\tau}_{\text{t-test}}$, provided that $\hat{m}_{i}$ is a good approximation of $m_{i}$, which is satified when $\hat{t}_{i}$ and $\hat{c}_{i}$ closely approximates $t_{i}$ and $c_{i}$. The variances can be calculated as follows:
@@
When it comes to methods for imputing potential outcomes, simply ignoring covariates and taking the mean of the observed outcomes in each treatment group makes the LOOP estimator equivalent to the student’s t-test estimator. However, the LOOP estimator outperforms $\hat{\tau}_{\text{t-test}}$ as long as the imputation of potential outcomes improves beyond the baseline mean imputation approach. In the work by Wu and Gagnon-Bartsch [Wu and Gagnon-Bartsch (2018)], Random Forest (RF) is recommended for imputing $t_{i}$ and $c_{i}$. This is because RF, as an enhancement over individual decision trees, not only improves accuracy, but also provides automatic variable selection, resulting in more computational efficiency.
In this paper, we calculate the t-test estimator $\hat{\tau}_{\text{t-test}}$, the Regression estimator $\hat{\tau}_{\text{Reg}}$, and the LOOP estimator $\hat{\tau}_{\text{LOOP}}$ to demonstrate that the LOOP estimator is an unbiased estimator of the ATE and achieves greater precision by adjusting for baseline covariates shown in equation 10. Furthermore, when using the LOOP estimator with RF, we can identify non-parametric relationships within the data that the Regression estimator fails to capture.